OpenStudy (anonymous):

A train leaves from point A for point B at 80 mph. After half an hour,another train leaves from point B for point A at 60 mph. which of the trains will be further from point A when they meet ? Math brain teaser

OpenStudy (anonymous):

ok I will help you if you help me

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

. There are 32 students in Miss Smith’s class. Eighteen are girls. What percent of the class are girls

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ok and it is ?

OpenStudy (anonymous):

\[\frac{18 girls }{32 students }\times 100\]

OpenStudy (anonymous):

if you can solve this then you can found the percent

OpenStudy (anonymous):

ok and how do I do that

OpenStudy (anonymous):

I got \[56 \frac{ 1 }{ 4 } \]

OpenStudy (anonymous):

ok good that is correct

OpenStudy (anonymous):

Ok cool

OpenStudy (anonymous):

so yours is 4800

OpenStudy (anonymous):

Right you solved the brain teaser

OpenStudy (anonymous):

relly thanks can I get a medle

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your very welcome